OpenStudy (poohdessey):

(h) The height of one solid limestone square pyramid is 24 m. A similar solid limestone square pyramid has a height of 30 m. The volume of the larger pyramid is 16,000 m3. Determine each of the following, showing all your work and reasoning. (a) The scale factor of the smaller pyramid to the larger pyramid in simplest form (b) The ratio of the area of the base of the smaller pyramid to the larger pyramid (c) Ratio of the volume of the smaller pyramid to the larger (i) The volume of the smaller pyramid

OpenStudy (poohdessey):

Please Help

OpenStudy (poohdessey):

@soprano.h.d0816

OpenStudy (poohdessey):

@izze17

OpenStudy (poohdessey):

@Tennis5518

OpenStudy (anonymous):

sry idk

OpenStudy (poohdessey):

@izze17 its ok

OpenStudy (poohdessey):

@Yuliacosta111

OpenStudy (yuliacosta111):

sorry idk

jhonyy9 (jhonyy9):

can you come back for help - i can give you

OpenStudy (poohdessey):

@jhonyy huh?

OpenStudy (error1603):

have you tried anything yet?

OpenStudy (math_man21):

ok i see

OpenStudy (mathstudent55):

You need to find the scale factor. You know that a side measuring 24 m corresponds to a side measuring 30 m. What is the scale factor of the side lengths?

OpenStudy (poohdessey):

@Error1603 yes i have.. and i need to learn how to do it, i tried asking my teacher last week she never got back to me

OpenStudy (math_man21):

its not a or d

OpenStudy (poohdessey):

@mathstudent55 thescale factor is 4/5

OpenStudy (math_man21):

should say i

OpenStudy (poohdessey):

@math_man21 its not multiple choiceis free response .

OpenStudy (mathstudent55):

Great. Since 24/30 = 4/5 = 0.8 That is correct. You have answered part (a)

OpenStudy (math_man21):

oh i see what we are doing now

OpenStudy (mathstudent55):

Now let's look at part (b). We are dealing with two square prisms. That means the bases are squares. Let's say the base of the larger prism measures L by L. Since the scale factor is 0.8, what does the base of the smaller prism measure?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Since the scale factor is 0.8, and the length and width of the larger prism are both L, what are the length and width of the base of the smaller prism?

OpenStudy (poohdessey):

@mathstudent55 so how would I find the base of the smaller prism? do I ned to divide 16,00 by 30

OpenStudy (mathstudent55):

No. Please just answer my question. The prisms have a scale factor of 0.8 Every linear dimension in the smaller prism is 0.8 the size of the corresponding dimension in the larger prism. If a measure is L in the larger prism, what is it in the smaller prism? (Hint: use L and the scale factor.)

OpenStudy (mathstudent55):

Hint 2: We know the heights are 30 m and 24 m. What do you multiply 30 m by to get 24 m?

OpenStudy (poohdessey):

ohhh you would multilply 30m by 0.8 to get 24m

OpenStudy (poohdessey):

@mathstudent55

OpenStudy (mathstudent55):

Correct. If you multiply 30 m by 0.8 to get 24 m because 0.8 is the scale factor, you multiply every linear dimension by 0.8 also. If the dimensions of the base (length and width) are both L in the larger prism, then in the smaller prism the length and width are both 0.8L Ok?

OpenStudy (mathstudent55):

Now we can answer part (b). What is the ratio of the area of the smaller base to the larger base? Small base: square with side 0.8L Large base: square with side L You have the sides of two squares. Find their areas and divide the smaller area by the larger area. Leave the result as a reduced fraction since they want a ratio.

OpenStudy (poohdessey):

so in order for me to find the area i would multiply 24 times 0.8 for the small base and 30 times ..... what for the larger one

OpenStudy (poohdessey):

@mathstudent55

OpenStudy (mathstudent55):

No. 24 m is the height. It has nothing to do with the area of the base. The base of the larger prism is a square that measures L by L. Forget the rest of this problem for a moment. Here is a simple question. If the side of a square has length L, what is the area of the square?

OpenStudy (poohdessey):

L times L

OpenStudy (mathstudent55):

Correct. The area is L^2 The larger prism has a base shaped like a square. Its side measures L. The area of the base of the larger prism is L^2

OpenStudy (mathstudent55):

The smaller prism also has a square as its base. We already concluded above that the sides of the base of the smaller prism measure 0.8L What is the area of a square with a side of length 0.8L?

OpenStudy (poohdessey):

.64l times L

OpenStudy (mathstudent55):

\(A_{square} = s^2\) Square with side L: \(A = L^2\) Square with side 0.8L: \(A = (0.8L)^2 = 0.8^2 \times L^2 = 0.64L^2\)

OpenStudy (mathstudent55):

Right. The area of the small base is 0.64L^2 The area of the large base is L^2 Now we need the ratio of the small area tot eh large area: \(\dfrac{0.64L^2}{L^2}\) You can cancel the L^2 in the numerator and denominator. What is the ratio of the bases of the small prism to the large prism?

OpenStudy (poohdessey):

.64

OpenStudy (mathstudent55):

Correct. The ratio is 0.64 to 1 We can write it as: \(\dfrac{0.64}{1} = \dfrac{64}{100} = \dfrac{16}{25} \)

OpenStudy (mathstudent55):

Now notice this. The ratio of the sides is 4/5, remember that?

OpenStudy (poohdessey):

Yes

OpenStudy (mathstudent55):

What is the ratio of the areas? We just found out it is 16/25 Do you notice that the ratio of the areas is the square of the ratio of the sides? 16/25 is the square of 4/5

OpenStudy (poohdessey):

ohh ok I see

OpenStudy (mathstudent55):

Now we can do part (c) The volume of a prism is area of the base times height. Large Prism Base Area: L^2 Small prism Base area: 0.64L^2 Large prism height: 30 m Small prism height: 0.8(30 m) Volume: L^2 * 30 m Volume: 0.64L^2 * 0.8(30 m) Ratio of volumes (smaller to larger): \(\dfrac{0.64L^2 \times 0.8(30 ~m)}{L^2 \times 30 ~m} \) \(\dfrac{0.64\cancel{L^2} \times 0.8(\cancel{30 ~m})}{\cancel{L^2} \times \cancel{30 ~m}~~1} \) \(= \dfrac{0.64 \times 0.8}{1}\) \(=\dfrac{0.512}{1}\) \(=\dfrac{512}{1000} = \dfrac{64}{125} \) The ratio of the volumes is the cube of the scale factor. \((\dfrac{4}{5})^3 = \dfrac{64}{125} \)

OpenStudy (mathstudent55):

Since the ratio of the volumes is 0.512/1, just multiply the volume of the larger prism by 0.512 to find the volume of the smaller prism. That is what part (d) asks.

OpenStudy (mathstudent55):

Sorry, but gtg. Bye

OpenStudy (poohdessey):

@mathstudent55 thnks!

OpenStudy (mathstudent55):

You're welcome.